what is the value of x for this equation?
1/5(x+0.5)+5.24=3/2x+7/10(x+2.2)

Found 2 solutions by MathLover1, MathTherapy:
Answer by MathLover1(20849)
Answer by MathTherapy(10552)


what is the value of x for this equation?
1/5(x+0.5)+5.24=3/2x+7/10(x+2.2)

2(x + 0.5) + 10(5.24) = 5(3)(x) + 7(x + 2.2) ------ Multiplying by LCD, 10
2x + 1 + 52.4 = 15x + 7x + 15.4
2x + 53.4 = 22x + 15.4
2x - 22x = 15.4 - 53.4
- 20x = - 38